Maria is now four times as old as Kate. Four years ago, Maria was six times as old as Kate. Find each of their actual ages now.

Answer by hheuler(6)

Let x represent the age of Kate four years ago. At the same time, Maria was 6x years old.
At the present time, Kate will be x + 4 years old and Maria will be 6x + 4 years old. We are told that Maria is now four times as old as Kate. So we have the following equation to solve.
6x + 4 = 4(x + 4)
6x - 2x = 16 - 4
2x = 12
x = 6
So Kate is now 6 + 4 = 10 years old and Maria is now 6(6) + 4 = 40 years old.